Use technology to find the P-value for a two-tailed test with

n equals=15

and test statistic t equals negative t=−2.143.
Answer #1

We have given that, test statistic t = -2.143

n = 15

Hence, degrees of freedom = n - 1 = 15-1 = 14

The test is two tailed test.

P-value = 2 * P( T < test statistic)

P-value = 2 * P (T < -2.143)

Using technology that is Excel we use the command,

=TDIST(2.143, 14, 2) we get the value 0.0502.

Hence, P-value = 0.0502